Circular No. 8359

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004cj–2004cq Further to IAUC 8302, A. Connolly reports the discovery by the ‘SDSS

Collaboration’ of several supernovae, detected during its spectroscopic survey operations in 45-min-exposure spectra (range ≈ 380–920 nm) taken with the SDSS multi-fiber spectrograph (with a field diameter of 3′′). The positions given for the supernovae below are those of the fiber center; each supernova is at the center of its host galaxy, except for SN 2004ck, which is 7′′ east and 8′′ north of its host’s center. Each supernova was identified by subtracting the underlying galaxy spectrum and then fitting supernova spectral templates to estimate the age of the supernova. The r magnitudes (on the AB system; cf. Fukugita et al. 1996, A.J. 111, 1748) for each supernova are estimated after subtracting the galaxy component. Also given below is the type and age (time in days after maximum; typical uncertainty ± 5 days) for each supernova. Redshifts for the host galaxies (consistent with those measured for the supernovae): 2004cj, 0.102; 2004ck, 0.029; 2004cl, 0.112; 2004cm, 0.005; 2004cn, 0.024; 2004co, 0.028; 2004cp, 0.054; 2004cq in UGC 9882, 0.022.

SN  2004 UT  α_{2000}  δ_{2000}  r Type (Age)
2004cj Jan. 29.34 10^{h}22^{m}50.^{s}12 +11^{o}42′10′′.8 20.1 Ia (−3)
2004ck Feb. 26.17 10 14 35.20 +11 52 05.7 19.8 Ia (+69)
2004cl Mar. 20.11 8 29 33.47 + 8 52 05.4 21.5 Ia (−4)
2004cm Mar. 24.46 14 07 24.96 +55 06 11.0 18.8 II (+21)
2004cn Apr. 17.27 12 27 28.11 +42 20 28.6 21.1 Ia (+43)
2004co Apr. 17.38 13 28 34.01 +41 51 08.2 17.4 Ia (+7)
2004cp May 24.21 15 40 24.76 +32 51 57.2 19.6 Ia (+28)
2004cq May 24.26 15 31 19.76 +40 16 52.1 17.5 Ia (+27)

SUPERNOVA 2004cr IN UGC 11603 Further to IAUC 8357, M. Armstrong, reports his discovery of an ap-

parent supernova (mag 17.8) on unfiltered CCD images taken on June 17.035 and 18.965 UT with a 0.35-m reflector. SN 2004cr is located at α = 20^{h}36^{m}06^{s}.84, δ = +63^{o}44′12′′.4 (equinox 2000.0), which is ≈ 45′′.9 west and 10′′.6 south of the center of UGC 11603. Nothing was visible at this position on his images taken on 2001 Aug. 5, 31, and 2003 July 13 (limiting mag ∼ 19.5) or on Palomar Sky Survey red and blue plates.

2004 June 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green